The total value of nickels and dimes in a bank is $2.50. If the nickels were dimes and the dimes were nickels, the total value would have been $2.00. How many nickels are there in the bank?

Answer by richwmiller(17219)

5n+10d=250
10n+5d=200
20n+10d=400
5n+10d=250
15n=150
n=10
5n+10d=250
50+10d=250
10d=200
d=20
check
10*10+5*20=200
100+100=200
ok